# Deep Learning based Key Information Extraction from Business Documents: Systematic Literature Review

### v1

@@ 7. Research agenda
    Novel datasets: Only a small number of public datasets are commonly used. For example, more than half of the sequence-based approaches are pre-trained using (subsets of) IIT-CDIP. One problem with this dataset in particular is the relatively low image quality, which no longer meets today’s standards. These properties have a direct impact on the use of corresponding approaches in real-world settings, which typically include documents with better image quality.
@@
    In addition to pre-training datasets, efforts should also be made to construct novel benchmark datasets, ideally based on different document types compared to receipts or forms which are most commonly considered, as also discussed in (Li et al., 2023a; Skalický et al., 2022). Besides different domains, newly created datasets should also be designed to more closely resemble documents found in real-world scenarios. (Yang et al., 2023a; Zhang et al., 2023b; Wang et al., 2023b) have shown, that existing dominant datasets have numerous shortcomings in this respect. Also, common benchmarks such as SROIE and FUNSD show a high degree of layout replication in training and test partitions (Laatiri et al., 2023), which distorts the overall validity the reported evaluation results. One possible solution to construct novel datasets can be the generation of synthetic documents, however, there is a risk that synthetic documents do not have the properties of real-world documents in terms of layout variety and complexity (Skalický et al., 2022).
  2. (2)
@@
    Consistent evaluation: The analysis in section 5.3 has shown that the papers are very heterogeneous in terms of their evaluation setup. In addition, the authors often do not specify their evaluation methodology in much detail, which raises questions on how exactly the evaluation results were obtained. Therefore, a quantitative comparison of the approaches is not always meaningful. This is especially problematic since the understanding of improving the state of the art in this research area is often associated with an increase in performance on benchmark datasets compared to existing work. Some public competitions, such as SROIE, provide a dedicated evaluation protocol for a consistent ranking of methods.  [^10] The leaderboard for the KIE task can be found at https://rrc.cvc.uab.es/?ch=13&com=evaluation&task=3. However, this is the exception rather than the rule.
    Therefore, it would be desirable to agree on a consistent approach or implement a centralized evaluation toolkit that could be used for different benchmark datasets, where a specific set of metrics (e.g., Precision, Recall, F1) is implemented. (Borchmann et al., 2021) have proposed a reference implementation for a DU benchmark, but it is not yet widely used.
@@
    We also advocate a more frequent use of string-based evaluation metrics, as they allow for a better assessment in real-world applications. However, only one third of the analyzed manuscripts used such string-based evaluation setups. It is also surprising that only about a third of the papers present the results on field-level. In this regard, we advocate that such an evaluation should be presented more frequently in future work, as it gives a good indication of whether an approach faces problems with certain fields, which in turn allows for a more in-depth analysis. Last, but not least, authors should adopt entity-level performance measures, as it is more suitable to assess the performance for practical applications where the extracted entities are processed as a whole (Zhang et al., 2023b).
  3. (3)
@@
    Pre-processing methods: There exists a dedicated research area that investigates techniques for pre-processing document images. An example of a sophisticated approach is proposed by (Saifullah et al., 2023b). In this context, synergy effects with this research area should be exploited, in particular the integration of corresponding techniques into KIE systems, as also discussed by (Baviskar et al., 2021). Related work often does not include pre-processing techniques, or, if they are used, only simple methods such as Deskewing. During the life cycle of a document in real-world settings, there are various points at which image quality can be impaired, for example due to introduced scan artifacts by digitization steps or as a consequence of multiple compression and transmission procedures (Alaei et al., 2024). (Yang et al., 2024) show, that quality aspects such as image noise or fonts can have significant impacts on the performance of DU systems. Therefore, more emphasis should be put on the adoption of respective methods in DU pipelines.
  4. (4)
@@
    Tokenization: In particular sequence-based KIE approaches utilize tokenizers for the (textual) encoding of input documents into token sequences, which are subsequently supplemented by visual and positional embeddings. Whenever the KIE systems incorporate pre-trained language models such as BERT as their encoder backbone, they usually also adopt the corresponding tokenizer without further adjustments. However, there have been studies highlighting the challenges associated with the adoption of unaltered tokenizers to different domains (Nayak et al., 2020). To this end, several methods have been proposed to adjust tokenizers to new domains, which improves the performance on downstream tasks. Examples are domain-specific augmentations of the original tokenizers’ vocabulary (Tai et al., 2020; Sachidananda et al., 2021), or a careful investigation of training data, pre-tokenization setups and other changes to the vocabulary (Dagan et al., 2024).
@@
  5. (5)
    Generalizability: (He et al., 2023a) have shown that existing KIE methods lack adequate generalization capabilities, which may be due to the limitations associated with the datasets as mentioned before. Furthermore, the aspect of generalizability is often not considered in detail. Instead, it is usually only implicitly considered when the approaches are evaluated based on multiple benchmark datasets with different document types.
@@
    Domain knowledge: As discussed in chapter 6.2, only a few authors integrate domain knowledge into the proposed approaches and/or adopt a practice-oriented view in general. However, it seems promising to explicitly integrate existing domain knowledge into KIE systems, which is also shown in (Cui et al., 2021), as it contains valuable information about the corresponding business processes, their documents and how they need to be understood as a whole. The design of novel pre-training tasks could be one possibility for integrating these aspects.
  7. (7)
@@
    Real-world usability: The relevance of automated document processing in real-world scenarios is indisputable (see also the review by (Martínez-Rojas et al., 2023)). Because of this great importance, more efforts should be made to improve the overall practicability of KIE approaches. The previously mentioned research directions already address this matter. For example, one could consider pre-processing techniques to remove document image artifacts that are specific to real-world settings, such as removing stamps (Yang et al., 2023b). The aforementioned focus on generalization capabilities can also have a positive impact on the practicality of KIE systems since deploying a system that can properly process many different types of documents simultaneously could result in lower costs and less maintenance.
@@
    On top of that, additional aspects could be considered. For example, (Kivimäki et al., 2023) propose a method to provide confidence estimates for the extracted data. The authors emphasise that in industrial settings, the primary goal is typically to make decisions based on model predictions rather than the raw extracted data. Research in this direction could greatly assist the collaboration between KIE systems and human administrators during document processing tasks, as it is essential to validate automatically extracted data in real-world settings (Houy et al., 2019). (Sassioui et al., 2023) also emphasize the importance of real-time KIE in industry applications. In this regard, a focus on lightweight KIE solutions should be considered. This could for example include investigating the trade-off between model size and extraction performance more closely. (Hamri et al., 2023) have shown that a relatively small model with around 50 thousand parameters can produce competitive results compared to systems that consist of several hundred million parameters.
  8. (8)
@@
    End-to-end performance: The evaluation has shown that end-to-end approaches usually achieve inferior extraction results compared to systems that use an external OCR engine. In this regard, more research should be conducted on how to improve the text recognition step, as the end-to-end systems have a high potential due to their independence of OCR engines, also highlighted by (Sassioui et al., 2023). End-to-end approaches have received less attention so far, which is also reflected in the number of identified approaches. Nevertheless, there is a relatively strong increase in corresponding methods, especially since 2022.
@@ 8. Conclusion
Research in the area of Key Information Extraction has seen an increased interest in recent years, mainly due to major advances in the field of Deep Learning. Nowadays, even visually-rich business documents with very complex layouts and high information-richness can be automatically processed by corresponding systems. To this end, this manuscript represents a systematic literature review covering the research on Key Information Extraction between 2017 and 2023 including 96 proposed approaches, with the aim of identifying the state of the art in this field and identifying potentials for further research. The identified methods were compared both qualitatively and quantitatively based on various characteristics.
@@
The analysis has shown that related work tends to follow a very narrow corridor in which already proven concepts are being successively refined and improved. In general, the approaches follow three key paradigms for representing document images, namely as sequences, graphs or grids. In detail, the approaches differ in their choice of architectures used for encoding and decoding the input documents, although it can also be seen that certain models are used more frequently (e.g., BERT-based models for textual inputs). Besides, novel concepts have been explored over time, such as OCR-independent and autoregressive methods, which on the one hand do not require an external OCR engine for preliminary text reading steps, and on the other hand can output arbitrary text, making them more flexible in terms of the downstream tasks they support. The authors investigate how different input modalities implicitly and explicitly contained in complex documents can be optimally integrated into the model architectures. In particular, visual cues that can be obtained from document images are increasingly incorporated into the models in order to improve their performance for more complex use cases. Much effort is also put into learning a general document understanding through deep learning models, which can then be used for Key Information Extraction. This is usually done by extensive and innovative pre-training tasks followed by specific fine-tuning steps. Another general observation is that the complexity of the corresponding models in terms of the number of parameters does not play a significant role and that even lightweight models can achieve promising extraction results.
@@
The research area is strongly characterized by the fact that an improvement of the state of the art is achieved by obtaining better extraction results on established benchmark datasets. However, a quantitative comparison of the presented results is not always meaningful, since the evaluation setups used by the authors are very heterogeneous. Furthermore, for most of the benchmarks, very good results have already been achieved (F1-scores above 0.97). The research area should therefore move away from focusing on improving benchmarks results and instead investigate innovations along other dimensions. Future work could focus on even more lightweight models in order to improve their practical applicability. It could also be investigated how the models can be designed in order to require less data for training procedures. We identified several other starting points for follow-up research based on the findings. These include proposing novel and more diverse datasets as well as consistent evaluation setups that allow for an adequate quantitative comparison. We also advocate that more attention should be paid to the real-world usability of corresponding approaches and to the integration of domain knowledge, since on the one hand document processing tasks play a key role in daily business workloads and, on the other hand, offer a special perspective on Key Information Extraction with unique requirements, but also possibilities.
@@ A. Business process perspective
B. Database search strings
@@
ACL intitle:((information AND extract*) OR ”entity extraction” OR ”entity recognition” OR document* OR ”visually rich”) AND (intitle:(”visually?rich” OR modal* OR image* OR invoice* OR receipt*) OR abstract:(”visually?rich” OR modal* OR image* OR invoice* OR receipt*) OR keywords:(”visually?rich” OR modal* OR image* OR invoice* OR receipt*)) 90
ACM Title:((information AND extract*) OR ”entity extraction” OR ”entity recognition” OR document* OR ”visually?rich”) AND ( Title:(”visually?rich” OR modal* OR image OR invoice OR receipt) OR Abstract:(”visually?rich” OR modal* OR image OR invoice OR receipt) OR Keyword:(”visually?rich” OR modal* OR image OR invoice OR receipt) ) AND NOT (Title:(handwrit*) OR Title:(histor*) OR Title:(table) OR Title:(web)) 248
AIS ((title:information AND title:extract*) OR title:”entity extraction” OR title:”entity recognition” OR title:document* OR title:”visually rich”) AND (”deep learning” OR ”artificial intelligence” OR ”neural network*”) &start_date=01/01/2017&end_date=12/31/2023 28
IEEE (((”Document Title”:information AND ”Document Title”:extract*) OR ”Document Title”:”entity extraction” OR ”Document Title”:”entity recognition” OR ”Document Title”:document* OR ”Document Title”:”visually rich”) AND ( (”Document Title”:”visually rich” OR ”Document Title”:modal* OR ”Document Title”:invoice OR ”Document Title”:receipt) OR (”Abstract”:”visually rich” OR ”Abstract”:modal* OR ”Abstract”:invoice OR ”Abstract”:receipt) OR (”Author Keywords”:”visually rich” OR ”Author Keywords”:modal* OR ”Author Keywords”:invoice OR ”Author Keywords”:receipt) ) NOT (”Document Title”:handwrit* OR ”Document Title”:histor* OR ”Document Title”:table)) 98
SD Year: 2017-2023; Title, abstract, keywords: ”visually rich” OR modal OR image OR invoice OR receipt; Title: (information AND extract) OR document OR ”visually rich” ; ”deep learning” OR ”artificial intelligence” OR cognitive OR intelligent OR ”neural network” 163
SL query=”visually+rich”+OR+modal*+OR+image+OR+invoice+OR+receipt & dc.title=information+extract* & date-facet-mode=between&facet-start-year=2017 & facet-end-year=2023 269
Total 896
@@ C. Numbering of the approaches
The manuscripts are sorted chronologically by year of publication. Within a year, they are sorted alphabetically by author name.

### v2

@@ 7. Research agenda
Novel datasets. Only a small number of public datasets are commonly used. For example, more than half of the sequence-based approaches are pre-trained using (subsets of) IIT-CDIP. One problem with this dataset in particular is the relatively low image quality, which no longer meets today’s standards. These properties have a direct impact on the use of corresponding approaches in real-world settings, which typically involve documents with better image quality.
@@
In addition to pre-training datasets, efforts should also be made to construct novel benchmark datasets, ideally based on different document types compared to commonly considered receipts – as also discussed in (Li et al., 2023a; Skalický et al., 2022). Besides different domains, newly created datasets should also be designed to more closely resemble documents found in real-world scenarios. (Ye et al., 2023; Zhang et al., 2023b; Wang et al., 2023b) have shown that existing dominant datasets have numerous shortcomings in this respect. Also, some datasets contain erroneous annotations and inconsistent labeling behavior (Lin et al., 2024). Furthermore, common benchmarks such as SROIE and FUNSD show a high degree of layout replication in training and test partitions (Laatiri et al., 2023), which distorts the overall validity the reported evaluation results. One possible solution to construct novel datasets can be the generation of synthetic documents, however, there is a risk that synthetic documents do not have the properties of real-world documents in terms of layout variety and complexity (Skalický et al., 2022).
@@
Consistent evaluation. The analysis in section 5.3 has shown that the papers are very heterogeneous in terms of their evaluation setup. In addition, the authors often do not specify their evaluation methodology in much detail, which raises questions on how exactly the evaluation results were obtained. Therefore, a quantitative comparison of the approaches is not always meaningful. This is especially problematic since the understanding of improving the state of the art in this research area is often associated with an increase in performance on benchmark datasets compared to existing work. Some public competitions, such as SROIE, provide a dedicated evaluation protocol for a consistent ranking of methods.  [^9] The leaderboard for the KIE task can be found at https://rrc.cvc.uab.es/?ch=13&com=evaluation&task=3. However, this is the exception rather than the rule.
@@
Therefore, it would be desirable to agree on a consistent approach or implement a centralized evaluation toolkit that could be used for different benchmark datasets, where a specific set of metrics (e.g., Precision, Recall, F1) is implemented. (Borchmann et al., 2021) have proposed a reference implementation for a DU benchmark, but it is not yet widely used. We also advocate a more frequent use of string-based evaluation metrics, as they allow for a better assessment in real-world applications. However, only one third of the analyzed manuscripts used such string-based evaluation setups. It is also surprising that only about a third of the papers present the results on field-level. In this regard, we advocate that such an evaluation should be presented more frequently in future work, as it gives a good indication of whether an approach faces problems with certain fields, which in turn allows for a more in-depth analysis, as also discussed by (Nourbakhsh et al., 2024). Last, but not least, authors should adopt entity-level performance measures, as it is more suitable to assess the performance for practical applications where the extracted entities are processed as a whole (Zhang et al., 2023b). The need for performance measures that provide a better indication of real-world performance is also addressed by (Khang et al., 2025), which propose a novel metric that shall mitigate the aforementioned shortcomings.
@@
Pre-processing methods. There exists a dedicated research area that investigates techniques for pre-processing document images. An example of a sophisticated approach is proposed by (Saifullah et al., 2023b). In this context, synergy effects with this research area should be exploited, in particular the integration of corresponding techniques into KIE systems (Baviskar et al., 2021). Related work often does not include pre-processing techniques, or, if they are used, only simple methods such as Deskewing. During the life cycle of a document in real-world settings, there are various points at which image quality can be impaired, for example due to introduced scan artifacts by digitization steps or as a consequence of multiple compression and transmission procedures (Alaei et al., 2023). (Yang et al., 2024a) show, that quality aspects such as image noise or fonts can have significant impacts on the performance of DU systems. Therefore, more emphasis should be put on the adoption of respective methods in DU pipelines.
@@
Tokenization. In particular sequence-based KIE approaches utilize tokenizers for the encoding of input documents into token sequences, which are subsequently supplemented by visual and positional embeddings. Whenever the KIE systems incorporate pre-trained language models such as BERT as their encoder backbone, they usually also adopt the corresponding tokenizer without further adjustments. However, there have been studies highlighting the challenges associated with the adoption of unaltered tokenizers to different domains (Nayak et al., 2020). To this end, several methods have been proposed to adjust tokenizers to new domains, which improves the performance on downstream tasks. Examples are domain-specific augmentations of the original tokenizers’ vocabulary (Tai et al., 2020; Sachidananda et al., 2021), or a careful investigation of training data, pre-tokenization setups and other changes to the vocabulary (Dagan et al., 2024).
@@
Generalizability. (He et al., 2023a) have shown that existing KIE methods lack adequate generalization capabilities, which may be due to the limitations associated with the datasets as mentioned before. Furthermore, the aspect of generalizability is often not considered in detail. Instead, it is usually only implicitly considered when the approaches are evaluated based on multiple benchmark datasets with different document types.
@@
Domain knowledge. As discussed in chapter 6.3, only a few authors integrate domain knowledge into the proposed approaches or adopt a practice-oriented view in general. However, it seems promising to integrate existing domain knowledge into KIE systems, which is also discussed in (Cui et al., 2021), as it contains valuable information about the corresponding business processes, their documents and how they need to be understood as a whole. The design of novel pre-training tasks could be one possibility for integrating these aspects.
@@
Real-world usability. The relevance of automated document processing in real-world scenarios is indisputable (see also the review by (Martínez-Rojas et al., 2023)). Because of this great importance, more efforts should be made to improve the overall practicability of KIE approaches. The previously mentioned research directions already address this matter. For example, one could consider pre-processing techniques to remove document image artifacts that are specific to real-world settings, such as removing stamps (Yang et al., 2023b). The aforementioned focus on generalization capabilities can also have a positive impact on the practicality of KIE systems since deploying a system that can properly process many different types of documents simultaneously could result in lower costs and less maintenance.
@@
On top of that, additional aspects could be considered. For example, (Kivimäki et al., 2023) propose a method to provide confidence estimates for the extracted data. The authors emphasize that in industrial settings, the primary goal is typically to make decisions based on model predictions rather than the raw extracted data. Research in this direction could greatly assist the collaboration between KIE systems and human administrators during document processing tasks, as it is essential to validate automatically extracted data in real-world settings (Houy et al., 2019). (Sassioui et al., 2023) also emphasize the importance of real-time KIE in industry applications. In this regard, a focus on lightweight KIE solutions should be considered. This could for example include investigating the trade-off between model size and extraction performance more closely. (He et al., 2023b) have shown that a relatively small model with around 50 thousand parameters can produce competitive results compared to systems that consist of several hundred million parameters.
@@
End-to-end performance. The evaluation has shown that end-to-end approaches usually achieve inferior extraction results compared to systems that use an external OCR engine. In this regard, more research should be conducted on how to improve the text recognition step, as the end-to-end systems have a high potential due to their independence of OCR engines, also highlighted by (Sassioui et al., 2023). End-to-end approaches have received less attention so far, which is also reflected in the number of identified approaches. Nevertheless, there is a relatively strong increase in corresponding methods, especially since 2022.
@@ 8. Conclusion
Research in the area of Key Information Extraction has seen an increased interest in recent years, mainly due to major advances in the field of Deep Learning. Nowadays, even visually-rich business documents with very complex layouts and high information-richness can be automatically processed by corresponding systems. To this end, this manuscript represents a systematic literature review covering the research on Key Information Extraction between 2017 and 2024 including 130 proposed approaches, with the aim of identifying the state of the art in this field and identifying potentials for further research. The identified methods were compared both qualitatively and quantitatively based on various characteristics.
@@
The analysis has shown that related work tends to follow a very narrow corridor in which already proven concepts are being successively refined and improved. In general, the approaches follow the same paradigms for representing document images, namely as sequences, graphs or grids. In detail, the approaches differ in their choice of architectures used for encoding and decoding the input documents, although it can also be seen that certain models are used more frequently (e.g., BERT-based models for textual inputs). Besides, novel concepts have been explored over time, such as OCR-independent and autoregressive methods, which on the one hand do not require an external OCR engine for preliminary text reading steps, and on the other hand can output arbitrary text, making them more flexible in terms of the downstream tasks they support. The authors investigate how different input modalities implicitly and explicitly contained in complex documents can be optimally integrated into the model architectures. In particular, visual cues from document images are increasingly incorporated into the models in order to improve their performance for more complex use cases. Much effort is also put into learning a general document understanding through Deep Learning models, which can then be used for Key Information Extraction. This is usually done by extensive and innovative pre-training tasks followed by specific fine-tuning steps. Another general observation is that the complexity of the corresponding models in terms of the number of parameters does not play a significant role and that even lightweight models can achieve promising extraction results.
@@
The research area is strongly characterized by the fact that an improvement of the state of the art is achieved by obtaining better extraction results on established benchmark datasets. However, a quantitative comparison of the presented results is not always meaningful, since the evaluation setups used by the authors are very heterogeneous. Furthermore, for most of the benchmarks, very good results have already been achieved (F1-scores above 0.97). The research area should therefore move away from focusing on improving benchmarks results and instead investigate innovations along other dimensions. Future work could focus on even more lightweight models in order to improve their practical applicability. It could also be investigated how the models can be designed in order to require less data for training procedures. We identified several other starting points for follow-up research based on the findings. These include proposing novel and more diverse datasets as well as consistent evaluation setups that allow for an adequate quantitative comparison. We also advocate that more attention should be paid to the real-world usability of corresponding approaches. This includes the integration of domain knowledge, since on the one hand document processing tasks play a key role in daily business workloads and, on the other hand, offer a special perspective on Key Information Extraction with unique requirements, but also possibilities.
@@ A. Business process perspective
B. Details for search procedure
B.1. Database search strings
@@
ACL intitle:((information AND extract*) OR ”entity extraction” OR ”entity recognition” OR document* OR ”visually rich”) AND (intitle:(”visually?rich” OR modal* OR image* OR invoice* OR receipt*) OR abstract:(”visually?rich” OR modal* OR image* OR invoice* OR receipt*) OR keywords:(”visually?rich” OR modal* OR image* OR invoice* OR receipt*)) 100
ACM Title:((information AND extract*) OR ”entity extraction” OR ”entity recognition” OR document* OR ”visually?rich”) AND ( Title:(”visually?rich” OR modal* OR image OR invoice OR receipt) OR Abstract:(”visually?rich” OR modal* OR image OR invoice OR receipt) OR Keyword:(”visually?rich” OR modal* OR image OR invoice OR receipt) ) AND NOT (Title:(handwrit*) OR Title:(histor*) OR Title:(table) OR Title:(web)) 300
AIS ((title:information AND title:extract*) OR title:”entity extraction” OR title:”entity recognition” OR title:document* OR title:”visually rich”) AND (”deep learning” OR ”artificial intelligence” OR ”neural network*”) &start_date=01/01/2017&end_date=12/31/2023 35
IEEE (((”Document Title”:information AND ”Document Title”:extract*) OR ”Document Title”:”entity extraction” OR ”Document Title”:”entity recognition” OR ”Document Title”:document* OR ”Document Title”:”visually rich”) AND ( (”Document Title”:”visually rich” OR ”Document Title”:modal* OR ”Document Title”:invoice OR ”Document Title”:receipt) OR (”Abstract”:”visually rich” OR ”Abstract”:modal* OR ”Abstract”:invoice OR ”Abstract”:receipt) OR (”Author Keywords”:”visually rich” OR ”Author Keywords”:modal* OR ”Author Keywords”:invoice OR ”Author Keywords”:receipt) ) NOT (”Document Title”:handwrit* OR ”Document Title”:histor* OR ”Document Title”:table)) 152
SD Year: 2017-2023; Title, abstract, keywords: ”visually rich” OR modal OR image OR invoice OR receipt; Title: (information AND extract) OR document OR ”visually rich” ; ”deep learning” OR ”artificial intelligence” OR cognitive OR intelligent OR ”neural network” 198
SL query=”visually+rich”+OR+modal*+OR+image+OR+invoice+OR+receipt & dc.title=information+extract* & date-facet-mode=between&facet-start-year=2017 & facet-end-year=2023 348
Total 1133
B.2. Prisma diagram
Figure SF2. PRISMA flow diagram
[…]
Literature search procedure according to PRISMA flow diagram
@@ C. Numbering of the approaches
The manuscripts are sorted chronologically by year of publication. Within a year, they are sorted alphabetically by author names.
@@
92 (Ye et al., 2023) (2023) UReader: Universal OCR-free Visually-situated Language Understanding with Multimodal Large Language Model
@@
97 (Ali and Roy, 2024) (2024) Enhancing Document Information Analysis with Multi-Task Pre-training: A Robust Approach for Information Extraction in Visually-Rich Documents
98 (Bai et al., 2024) (2024) AIE-KB: Information Extraction Technology with Knowledge Base for Chinese Archival Scenario
99 (Biescas et al., 2024) (2024) GeoContrastNet: Contrastive Key-Value Edge Learning for Language-Agnostic Document Understanding
100 (Dai et al., 2024) (2024) GraphMLLM: A Graph-Based Multi-level Layout Language-Independent Model for Document Understanding
101 (Feng et al., 2024) (2024) DocPedia: unleashing the power of large multimodal model in the frequency domain for versatile document understanding
102 (Fujitake, 2024) (2024) LayoutLLM: Large Language Model Instruction Tuning for Visually Rich Document Understanding
103 (Gbada et al., 2024a) (2024a) Information Extraction from Visually Rich Documents Using Directed Weighted Graph Neural Network
104 (Gbada et al., 2024b) (2024b) Multimodal weighted graph representation for information extraction from visually rich documents
106 (Hu et al., 2024b) (2024b) mPLUG-DocOwl 1.5: Unified Structure Learning for OCR-free Document Understanding
107 (Hu et al., 2024a) (2024a) UniVIE: A Unified Label Space Approach to Visual Information Extraction from Form-Like Documents
108 (Huang et al., 2024) (2024) MCKIE: Multi-class Key Information Extraction from Complex Documents Based on Graph Convolutional Network
109 (Kubade et al., 2024) (2024) Reformulating Key-Information Extraction as Next Sentence Prediction for Hierarchical Data
110 (Kumar et al., 2024) (2024) ProtoNER: Few Shot Incremental Learning for Named Entity Recognition Using Prototypical Networks
111 (Li et al., 2024) (2024) DocPointer: A parameter-efficient Pointer Network for Key Information Extraction
112 (Lin et al., 2024) (2024) PEneo: Unifying Line Extraction, Line Grouping, and Entity Linking for End-to-end Document Pair Extraction
113 (Liu et al., 2024) (2024) HRVDA: High-Resolution Visual Document Assistant
114 (Luo et al., 2024a) (2024a) LayoutLLM: Layout Instruction Tuning with Large Language Models for Document Understanding
115 (Luo et al., 2024b) (2024b) Incorporating multivariate semantic association graphs into multimodal networks for information extraction from documents
116 (Minouei et al., 2024) (2024) Embedding Layout in Text for Document Understanding Using Large Language Models
117 (Perot et al., 2024) (2024) LMDX: Language Model-based Document Information Extraction and Localization
118 (Qi et al., 2024) (2024) GVDIE: A Zero-Shot Generative Information Extraction Method for Visual Documents Based on Large Language Models
119 (Siyuan et al., 2024) (2024) LayoutPointer: A Spatial-Context Adaptive Pointer Network for Visual Information Extraction
120 (The et al., 2024) (2024) One-Shot Transformer-Based Framework for Visually-Rich Document Understanding
121 (Tran et al., 2024) (2024) Efficiency evaluation of filter sizes on graph convolutional neural networks for information extraction from receipts
122 (Trivedi et al., 2024) (2024) GDP: Generic Document Pretraining to Improve Document Understanding
123 (Tu et al., 2024) (2024) UNER: A Unified Prediction Head for Named Entity Recognition in Visually-rich Documents
124 (Wan et al., 2024) (2024) OMNIPARSER: A Unified Framework for Text Spotting, Key Information Extraction and Table Recognition
125 (Wang and Wei, 2024) (2024) RobustLayoutLM: Leveraging Optimized Layout with Additional Modalities for Improved Document Understanding
126 (Wang et al., 2024b) (2024b) DocLLM: A Layout-Aware Generative Language Model for Multimodal Document Understanding
127 (Wang et al., 2024a) (2024a) LiLTv2: Language-substitutable Layout-Image Transformer for Visual Information Extraction
128 (Xu et al., 2024) (2024) EntityLayout: Entity-Level Pre-training Language Model for Semantic Entity Recognition and Relation Extraction
129 (Yang et al., 2024b) (2024b) Light-Weight Multi-modality Feature Fusion Network for Visually-Rich Document Understanding
130 (Yoon et al., 2024) (2024) Language, OCR, Form Independent (LOFI) pipeline for Industrial Document Information Extraction